# Kirchhoff Meets Johnson: In Pursuit of Unconditionally Secure Communication

### v2

@@ VI Clash Of the Titans: KLJN vs QKD
Two prevalent QKD attacks are intercept-resend (IR) and photon number splitting (PNS) attacks [52, 17]. In the former, Eve intercepts the quantum signal sent by Alice and measures it. Based on the measurement result, she prepares a new quantum signal in the same measured state to be sent to Bob. Due to quantum weirdness, Eve can only detect correct states with a probability of $1/2$. In return, the overall error probability between Alice and Bob becomes $1/4$, which is relatively high and can be easily noticed by them. However, in the more sophisticated PNS attack, Eve exploits the fact that Alice transmits multiple photons for each bit. Specifically, Alice’s coherent source emits photons according to the Poisson distribution with a typical average photon number of $0.1-0.8$ (Poisson parameter $\alpha$) for weak laser pulses [30]. Despite its rare occurrence probability, in this case, Eve employs a beamsplitter to take one of the photons and pass the rest to Bob. To circumvent this, Alice can employ a decoy-state-based protocol. In this protocol, to represent the signal and decoy states, Alice transmits quantum states with different mean photon numbers to confuse Eve [53]. While the literature contains a rich set of other attacks involving heavy quantum physics and numerous debates about the practical vulnerabilities of QKD systems, we conclude our discussion at this point. Table I presents a high-level comparison of KLJN and QKD systems considering their major key performance indicators (KPIs), such as key rate, range, complexity, cost, and BER. Here, we use high, medium, and low labels based on specific KPIs for each system. For instance, QKD outperforms KLJN regarding key rate, benefiting from not requiring a circuit loop between two users. Nevertheless, sophisticated attacks utilizing photon groups can limit the key rate of practical QKD systems. However, QKD and KLJN systems have different power scaling laws and the target range is critical in their comparison. We note that the complexity of KLJN can be as high as QKD when fully defended but its cost can be very low if integrated on a chip. In the same table, we also list the corresponding major protocols associated with these two physics-based key distribution systems. We conclude that both designs can offer unconditional security by exploiting the laws of physics under certain conditions. However, the choice between these two key distribution approaches should be made carefully, depending on the requirements of the target applications. Evaluating the performance of two concepts through computer simulations and practical experiments is also a promising future research direction. At this point, we also pose the following interesting question: can these two different key distribution concepts merged to create an ultimate key distribution system? To conclude, both KLJN and QKD systems rely on specific laws of physics to create secure keys; however, they might require additional defense mechanisms in practice to preserve unconditional security. If implemented at the chip level, the low cost of KLJN systems makes them a potential candidate for short-range critical applications in which system designers can create effective Kirchhoff loops. On the other hand, QKD systems require dedicated hardware, have a much higher cost, and can be suitable for more advanced and long-range key distribution applications.
@@ VII Noise is in the Air
Breaking the Kirchhoff loop of KLJN and performing IM for resistances over wireless channels has led wireless thermal noise modulation (TherMod), which is illustrated in Fig. 7. In this setup, incoming information bits at the transmitter determine the index of the selected resistor similar to the KLJN scheme. This selection allows the modulation of the power spectral density of the generated thermal noise waveform, as initially demonstrated in the early study of [4]. Despite the conceptual similarity between KLJN and TherMod, the latter cannot provide unconditionally secure communication due to the non-existence of a noise loop between transmitter and receiver. However, if implemented by background noises, a transmitter with extremely low power consumption (less than $1$ microwatt as backscatter communication (BackCom) systems) can be obtained, which also provides stealth communication capabilities. Another advantage of TherMod is that its transmitter does not rely on existing RF signals to operate as ambient BackCom systems [54]. The concept of TherMod was envisioned in the early work of [4] by two parabolic antennas driven by variable resistances (open, short, and 50 ohms). An experimental demonstration of this scheme has been recently given in [55], where due to extremely low thermal noise powers, a short range and line-of-sight (LOS)-dominated communication setup is considered with highly directional horn antennas. Nevertheless, by simply indexing resistances, the authors reported a maximum data rate of $26$ bits/second for short distances and a maximum range of $7.3$ meters. Although some early designs have been reported in the literature to convey information by random signal parameters [56, 57, 58, 59], such as variance, we have provided a general framework by using resistance IM to convey information and obtained comprehensive theoretical results depending on specific system parameters, such as useful and disruptive noise powers [3]. Recently, the noise modulation concept is also proposed to use externally generated noise signals and a number of noise-alike waveforms are created to convey information bits through wireless channels [60]. Similar to the KLJN system, the task of the TherMod receiver is to extract the information embedded in the variance of the noise signals generated at the transmitter. For this purpose, the receiver software-defined radio (SDR) device can take samples from the received waveform and obtain complex baseband samples for further processing. Here, the level of the additional noise the receiver chain introduces should be as low as possible to make a reliable bit detection. For instance, as shown in Fig. 7, calculating the variance estimate $(\hat{\sigma}_{s}^{2})$, a decision on the transmitted bit can be made by comparing it with a suitable threshold. In contrast to resistor indexing, in the scheme of noise loop modulation [61], by creating a full-duplex noise communication loop between the transmitter and the receiver, the noise itself is modulated to obtain a solid secrecy at the PHY level for both additive white Gaussian noise and multipath fading channels. In other words, this two-way loop ensures that there is always a mixture of the user bits and noises over the air, and a third node with different noise processes cannot reveal the transmitted information, no matter its computational power. The scheme of noise loop modulation appears as a promising candidate to use the noise signals themselves for unconditional security over wireless channels, differently from artificial noise-based PHY security systems that mainly aim to degrade the signal-to-noise ratio of potential eavesdroppers [28]. On the downside, the data rate of this scheme is relatively limited, and it requires a full-duplex communication loop to operate. Furthermore, the effect of changing bit values at both sides (transients) should be carefully examined from the perspective of preserving the secrecy.
@@ VIII Conclusions and Future Directions
In this review article, under the central theme of Kirchhoff meets Johnson, we have first introduced the fundamentals of thermal noise-driven communication and the KLJN scheme. Then, using the KLJN scheme as a reference point, we have presented a historical journey toward unconditionally secure communication and revisited the longstanding yet critical challenge in this field: secure key distribution. In pursuing unconditionally secure communication empowered by physics-based key distribution schemes, we also conducted a detailed review of QKD systems and provided concise comparisons with KLJN systems. Lastly, we have explored over-the-air thermal noise transmission concepts and evaluated their promising potential for low-power, stealth, and secure communications. Potential research directions for thermal noise-driven communication include the generalization of TherCom schemes for an arbitrary number of resistors and external sources, the exploration of practical issues such as wire resistances, sampling imperfections, and timing errors, the proposal of suitable coding schemes or effective bit mapping algorithms to improve the error performance further, and the integration of these concepts into legacy communication systems. Mainly, addressing the limited key distribution range and key rate of KLJN is a challenge that needs resolution for its widespread practical use. QKD systems’ very high cost and fragile structure also require careful assessment for their practical applications. In this context, alternative key distribution solutions such as KLJN can be evaluated in some instances to complement their quantum-based counterparts. However, the security of both designs should be meticulously examined under practical imperfections and constraints. Furthermore, development of unconditionally secure PHY security signaling techniques empowered by noise-alike signals, is an exciting future research direction. We ask again: Is it game over for codebreakers after their relentless efforts to outwit codemakers? In other words, with the advent of unconditionally secure systems, has the millennia-old battle between codebreakers and codemakers finally concluded? While these complex queries do not yield quick answers, the emergence of physics-based secure key distribution techniques and their potential for widespread adoption soon imply that codebreakers would encounter a formidable challenge when attempting to breach unconditionally secure communication systems.

### v3

@@ VII Noise is in the Air
Breaking the Kirchhoff loop of KLJN and performing IM for resistances over wireless channels has led wireless thermal noise modulation (TherMod), which is illustrated in Fig. 8. In this setup, incoming information bits at the transmitter determine the index of the selected resistor similar to the KLJN scheme. This selection allows the modulation of the power spectral density of the generated thermal noise waveform, as initially demonstrated in the early study of [4]. Despite the conceptual similarity between KLJN and TherMod, the latter cannot provide unconditionally secure communication due to the non-existence of a noise loop between transmitter and receiver. However, if implemented by background noises, a transmitter with extremely low power consumption (less than $1$ microwatt as backscatter communication (BackCom) systems) can be obtained, which also provides stealth communication capabilities. Another advantage of TherMod is that its transmitter does not rely on existing RF signals to operate as ambient BackCom systems [54].
The concept of TherMod was envisioned in the early work of [4] by two parabolic antennas driven by variable resistances (open, short, and 50 ohms). An experimental demonstration of this scheme has been recently given in [55], where due to extremely low thermal noise powers, a short range and line-of-sight (LOS)-dominated communication setup is considered with highly directional horn antennas. Nevertheless, by simply indexing resistances, the authors reported a maximum data rate of $26$ bits/second for short distances and a maximum range of $7.3$ meters. Although some early designs have been reported in the literature to convey information by random signal parameters [56, 57, 58, 59], such as variance, we have provided a general framework by using resistance IM to convey information and obtained comprehensive theoretical results depending on specific system parameters, such as useful and disruptive noise powers [3]. Recently, the noise modulation concept is also proposed to use externally generated noise signals and a number of noise-alike waveforms are created to convey information bits through wireless channels [60].
Similar to the KLJN system, the task of the TherMod receiver is to extract the information embedded in the variance of the noise signals generated at the transmitter. For this purpose, the receiver software-defined radio (SDR) device can take samples from the received waveform and obtain complex baseband samples for further processing. Here, the level of the additional noise the receiver chain introduces should be as low as possible to make a reliable bit detection. For instance, as shown in Fig. 8, calculating the variance estimate $(\hat{\sigma}_{s}^{2})$, a decision on the transmitted bit can be made by comparing it with a suitable threshold.
In contrast to resistor indexing, in the scheme of noise loop modulation [61], by creating a full-duplex noise communication loop between the transmitter and the receiver, the noise itself is modulated to obtain a solid secrecy at the PHY level for both additive white Gaussian noise and multipath fading channels. In other words, this two-way loop ensures that there is always a mixture of the user bits and noises over the air, and a third node with different noise processes cannot reveal the transmitted information, no matter its computational power. The scheme of noise loop modulation appears as a promising candidate to use the noise signals themselves for unconditional security over wireless channels, differently from artificial noise-based PHY security systems that mainly aim to degrade the signal-to-noise ratio of potential eavesdroppers [28]. On the downside, the data rate of this scheme is relatively limited, and it requires a full-duplex communication loop to operate. Furthermore, the effect of changing bit values at both sides (transients) should be carefully examined from the perspective of preserving the secrecy.
@@ VIII Conclusions and Future Directions
In this review article, under the central theme of Kirchhoff meets Johnson, we have first introduced the fundamentals of thermal noise-driven communication and the KLJN scheme. Then, using the KLJN scheme as a reference point, we have presented a historical journey toward unconditionally secure communication and revisited the longstanding yet critical challenge in this field: secure key distribution. In pursuing unconditionally secure communication empowered by physics-based key distribution schemes, we also conducted a detailed review of QKD systems and provided concise comparisons with KLJN systems. Lastly, we have explored over-the-air thermal noise transmission concepts and evaluated their promising potential for low-power, stealth, and secure communications.
The major advantages and weaknesses of KLJN-based systems are summarized as follows. First, the KLJN systems can be deployed at a much lower cost if implemented at the chip level and can secure short- to medium-range links effectively without requiring significant maintenance. Utilizing more advanced detectors with joint noise voltage and current samples, the error probability of KLJN systems can be reduced further, which might help increase the range and key rate. As a result, KLJN-based solutions shine out for potential applications such as securing communications within data centers and local networks. However, the cost of the network will increase to defend the system against various attacks fully. The requirement for an active Kirchoff loop also hinders the flexibility of KLJN-based systems to operate at challenging conditions and long ranges.
@@
Potential future research and development directions for unconditionally secure communication systems include the generalization of TherCom schemes for an arbitrary number of resistors and external sources, the exploration of practical issues such as wire resistances, sampling imperfections, and timing errors, the proposal of suitable coding schemes or effective bit mapping algorithms to improve the error performance further, and the integration of these concepts into legacy communication systems. Mainly, addressing the limited key distribution range and key rate of KLJN is a challenge that needs resolution for its widespread practical use. Furthermore, developing novel KLJN variants that can be used for two-way secure data transfer is an interesting research direction. In such a setup, KLJN communicators can be directly used to exchange secure data between two users, but the transmission needs to be unconditionally secure for all bit cases. QKD systems’ very high cost and fragile structure also require careful assessment for their practical applications. In this context, alternative key distribution solutions such as KLJN can be evaluated in certain cases to complement their quantum-based counterparts. At this point, merging these two key distribution concepts might also be an interesting problem in creating an ultimate unconditionally key distribution system. However, the security of both designs should be meticulously examined under practical imperfections and constraints. Furthermore, developing unconditionally secure PHY security signaling techniques empowered by noise-alike signals and providing a complete uncertainty for hidden eavesdroppers at the PHY layer is an exciting future research direction.
We ask again: Is it game over for codebreakers after their relentless efforts to outwit codemakers? In other words, with the advent of unconditionally secure systems, has the millennia-old battle between codebreakers and codemakers finally concluded? While these complex queries do not yield quick answers, the emergence of physics-based secure key distribution techniques and their potential for widespread adoption soon imply that codebreakers would encounter a formidable challenge when attempting to breach unconditionally secure communication systems.